OpenStudy (jenniferjuice):

Choose the correct product of (3x − 8)^2 9x^2 − 64 9x^2 + 64 9x^2 − 48x + 64 9x^2 + 48x + 64

OpenStudy (uri):

3*3=? 8*8=?

OpenStudy (jenniferjuice):

9 and 64

OpenStudy (anonymous):

so with what they said, you end up with.... 9 x^2-48 x+64

OpenStudy (anonymous):

u know formula (a-b)^2=(a^2)-(2*a*b)+(b^2) so( 3x-8)^2=((3x)^2)-(2*3x*8)+(8^2) =(9x^2)-48x+64 option C

OpenStudy (uri):

Wait it's the (a+b)^2 formula? My bad.. (3-8)^2=(3x)^2-2(3)(8)+(8)^2 9x^2-48+64

OpenStudy (anonymous):

Its okay :P

OpenStudy (jenniferjuice):

wait ....so it's 9x^2 − 48x + 64 ..?

OpenStudy (uri):

Yeah